### 29 Feb 2024 v1

# Let LLMs Take on the Latest Challenges ! A Chinese Dynamic Question Answering Benchmark

Abstract
How to better evaluate the capabilities of Large Language Models (LLMs) is the focal point and hot topic in current LLMs research. Previous work has noted that due to the extremely high cost of iterative updates of LLMs, they are often unable to answer the latest dynamic questions well. To promote the improvement of Chinese LLMs’ ability to answer dynamic questions, in this paper, we introduce CDQA, a Chinese Dynamic QA benchmark containing question-answer pairs related to the latest news on the Chinese Internet. We obtain high-quality data through a pipeline that combines humans and models, and carefully classify the samples according to the frequency of answer changes to facilitate a more fine-grained observation of LLMs’ capabilities. We have also evaluated and analyzed mainstream and advanced Chinese LLMs on CDQA. Extensive experiments and valuable insights suggest that our proposed CDQA is challenging and worthy of more further study. We believe that the benchmark we provide will become the key data resource for improving LLMs’ Chinese question-answering ability in the future.
1 Introduction
Due to the excellent emergence capabilities and unified task paradigm, Large Language Models (LLMs) are undoubtedly the more popular stars in the field of Natural Language Processing (NLP) or Artificial Intelligence Wei et al. (2022); Li et al. (2023); Shanahan (2024). To promote the improvement of LLMs capabilities, more and more researchers have invested in building various LLMs evaluation benchmarks Chang et al. (2023); Huang et al. (2023a). In the era of LLMs, high-quality evaluation benchmarks allow researchers to better understand the capabilities of LLMs, thereby stimulating further research on how to enhance LLMs.
Question answering is an important and long-standing topic in NLP Rajpurkar et al. (2016); Joshi et al. (2017); He et al. (2018). Especially for LLMs, QA tasks have almost become the indispensable basic task in LLMs research Pan et al. (2024). Various forms of QA benchmarks can be used to measure the capabilities of LLMs in different dimensions Adlakha et al. (2022); Bosselut et al. (2022); Rein et al. (2023); Huang et al. (2023b). Recently, the introduction of English FreshQA Vu et al. (2023) has attracted widespread attention. It challenges LLMs through questions with dynamically changing answers, aiming to test LLMs’ mastery of the latest factual knowledge. Obviously, being able to answer the latest questions determines to some extent whether LLMs can truly move towards large-scale daily applications. Urgently, we note that there is still no such benchmark in the Chinese community, although LLMs in the Chinese scenario still face the same challenges and dilemmas, as shown in Table 1.
Static ACL 主会每年举办几次?
Question How many times does the ACL
annual meeting take place each year?
GPT-4’s 一年一次。
Answer Once a year.
Dynamic 下一次 ACL 将在哪里举办?
Question Where will the next ACL be held?
GPT-4’s 我无法提供相关信息。
Answer I can’t provide the information.
Table 1: Examples of static and dynamic questions. The GPT-4 is on Feb 11, 2024.
To let LLMs in Chinese scenarios take on the latest challenges and empower them to answer dynamic questions, in this work, we present CDQA, a Chinese Dynamic QA benchmark. Specifically, we design a semi-automatic data production pipeline to construct our benchmark. In this pipeline, we first automatically generate a large number of raw queries with the help of two LLMs with different roles, one is to extract key entities from the latest Chinese news, and the other is to automatically generate question queries based on the extracted entities that will be as the corresponding answers. Then we ask the well-trained annotators to filter, rewrite, and classify the automatically generated question samples to ensure the quality of CDQA. Through such a semi-automatic data construction method with human participation, we obtain 1,339 question-answer pairs for CDQA, classified by how frequently their answers change (i.e., fast-changing, slow-changing, and never-changing). The purpose of classifying CDQA samples by the frequency of answer changes is to provide finer-grained evaluation for LLMs, facilitating researchers to better perceive the true performance of LLMs.
Based on our constructed CDQA, we select a series of widely used and advanced LLMs in the Chinese community for evaluation. Results show that GPT-4 still ranks at the top with searched results from search engines, surpassing at most nearly 10 F1-recall scores with less hallucination than the second-best model, i.e., Deepseek-67B-Chat, although Chinese-oriented LLMs tend to have more internal knowledge than OpenAI models. Besides, in-context learning and prompting methods like Chain-of-Thought generally increase performances with searched evidence but also elicit more hallucinations in LLMs. For search engines in the open-book scenario, Google consistently takes advantage over Bing for all baseline models, showing great strength as a retriever for LLMs.
In summary, our contributions could be summarized as follows:
  1. 1.
    We first introduce the idea of using dynamic questions to challenge Chinese LLMs, which provides a new direction for the development of LLMs in Chinese community.
  2. 2.
    We construct the high-quality CDQA benchmark composed of dynamic questions, which will become an important data resource for promoting the progress of Chinese LLMs.
  3. 3.
    Extensive experiments and detailed analyses based on CDQA provide valuable insights and discoveries, which are instructive for subsequent research about how to enhance LLMs to handle dynamic questions.
2 Chinese Dynamic Question Answering (CDQA)
2.1 Overview
Our dataset, CDQA, mainly originates from latest news in Chinese Internet from different areas such as finance, daily life, politics, technology and so on. Besides, there are also queries collected from Chinese labelers. They represent the information-seeking cases of Chinese people. The generation pipeline could be illustrated in Figure 1. The dataset currently consists of 1,339 questions covering a range of topics with evolving answers which are mostly extracted entities from the raw corpus scraped from Chinese Internet and it is being regularly updated.
Figure 1: Data Generation Pipeline for CDQA dataset. We first collect Chinese News from Internet and then extract entities from these news passages. Based on GPT-4, we generate synthetic queries from passages and corresponding entities. Manual annotation is conducted to verify the synthetic data and extra human-crafted queries, providing the verified queries, answers and supportive evidence links.
2.2 Data Collection
We collect CDQA dataset in two stage. The first stage is automatic generations with Entity Extraction and Doc2Query, for which we use SeqGPT Yu et al. (2023), and GPT-4 OpenAI (2023), which could give great amount of raw question-answering pairs as SeqGPT extracts entities from latest Chinese news and GPT-4 is prompted into generating corresponding questions. For GPT-4 prompts, we use few-shot prompting in generating diverse questions from entities. The second stage is manual labeling from crowd-sourced workers. The Chinese labelers not only filter questions which are answered with biases, ambiguities and obsolete [^1] The answer should be only supported with the knowledge after January 1st 2019 except for static knowledge, i.e., never-changing. knowledge but also annotate with tags, check the correctness and rewrite the question answer pairs to be more time-related and dynamic. At the very beginning, the labelers are shown with pre-annotation examples and annotation guides.
Tags
The tags are annotated for questions and answers. For questions, we have the same taxonomy as FreshQA Vu et al. (2023). The questions are categorized as fast-changing, slow-changing, and never-changing. For answers, we categorize these entities or short texts as person, location, time, event, artificial work, group, nature, quantity and other. Therefore, we could evaluate the models’ latest world knowledge from various perspectives. The taxonomy is illustrated in Appendix A.
Quality Control
After getting the synthetic queries, the human annotators could rewrite and calibrate the questions and answers to make the QA pairs correct, consistent and dynamic. For example, annotators are required to provide the supporting evidence URLs along with correct answers using search engines. This calibration process could solidify our answers with supplementary valid information and help us better iterate the dataset as the generation process in the previous stage is not well-evaluated with supportive documents, let alone the correctness. Moreover, in order to facilitate the periodic updates, we filter out the questions with more than one valid answer.
For inter-annotator agreement, we randomly sample 100 examples from synthetic question-answer pairs and annotations from two annotators in the same annotation vendor are measured by acceptance (whether the pair is accepted or discarded), question tags and answer types. The ground-truth labels are provided by authors. For each category, we calculate their Cohen Kappa scores McHugh (2012). From Table 2, the averaged score across all types of annotations are above 63.1, representing “substantial agreement” for our dataset annotations.
Acceptance Question Tags Answer Types
Ann1 v.s. Ann2 62.3 87.2 96.6
GT v.s. Ann1 79.6 59.1 100
GT v.s. Ann2 47.3 68.3 100
Avg 63.1 71.5 98.9
Table 2: Inter-annotator agreement for different annotation sections are calculated by Cohen Kappa scores. Ann1/2 represents Annotator1/2 respectively and GT represents Ground Truth. Our annotations could be considered as “substantial agreement” as the average scores are above 60.
2.3 Regular Updates
Our dataset is highly sensitive to time since the ground truth is evolving along the world development. Therefore, we commit to updating the dataset regularly and researchers are strongly encouraged to stay tuned with our latest version for evaluation. And the datasets are mainly calibrated with information from Chinese Internet.
2.4 Data Statistics
Due to limitations in automatic query generation by GPT-4 and SeqGPT from the first stage, our dataset has low retention rate in which only 44.6% synthetic data are accepted by human annotators. Among the accepted data, 53.1% of them still need further modifications because of improper questions or wrong answers. For question tags, we have relatively balanced distributions between fast-changing and slow-changing questions with fewer never-changing questions. For answer types, we have biased distributions as nearly 70% of entities extracted from passages lie in “person” and “group” categories. This is because most of entities in first stage by automatic generation are “person” and “group”. However, question tags and answer types could be changed or calibrated over time by re-annotation of the dataset. These distribution graphs and more details about our dataset are in Appendix B.
2.5 Evaluation
As our dataset is constructed from Internet, our evaluation setup is based on retrieval-augmented question answering with search engines while the evaluation methods are answer rate and F1-recall. Our experiments consist of two settings: close-book and open-book. Overall, our evaluation provides a comprehensive understanding of current LLMs in factuality, especially for evolving knowledge. Besides, due to the safety implementation for different LLMs from helpful and harmless responses in training data Bai et al. (2022), F1-recall only counts on questions with effective responses while answer rate is used in representing the ratio of answered questions to the total questions, which is a practical metric for the real world application of LLMs and could directly indicate the degree of hallucination in generated responses.
Evaluation Metrics
For F1-recall, we calculate the ratio of common tokens between model-generated responses and ground truth to the ground truth. Specifically, we first segment the generated text and golden text into token lists using word segmentation tools  [^2] https://github.com/fxsjy/jieba, then calculate the ratio of tokens generated by models belonging to the golden token list to golden tokens. For answer rate, we directly calculate the ratio of effectively answered questions to total questions, i.e., responses of refusal are filtered out in our evaluation.
3 Experiments
Figure 2: Our prompts are formulated under this framework. Different prompting methods are used with different instructions $\mathbf{i}$. The Chinese version is in Appendix C.
Model fast-changing slow-changing never-changing
+0 +5 +10 +0 +5 +10 +0 +5 +10
Intermlm-20B-Chat 16.1 [16] 49.7 [0] 50.5 [0] 19.8 [16] 61.2 [16] 61.8 [0] 34.0 [16] 68.1 [16] 71.7 [16]
Aquila2-34B-Chat 14.9 [16] 49.4 (99.6%) [0] 51.5 [0] 17.7 [16] 60.4 [0] 62.5 [0] 35.6 [5] 69.2 [0] 69.4 [0]
Yi-34B-Chat 22.9 [16] 54.9 [16] 56.5 [16] 30.8 [16] 67.0 [16] 68.8 [16] 46.9 [16] 76.0 [16] 76.9 [16]
Deepseek-67B-Chat 24.3 [0] 57.5 [16] 58.4 [16] 37.2 [16] 68.3 [16] 70.0 [16] 53.1 [0] 78.8 [16] 79.2 [5]
Qwen-72B-Chat 18.6 [16] 51.7 [16] 54.7 [16] 16.9 [16] 63.1 [0] 64.6 [16] 41.5 [16] 74.6 [16] 75.4 [16]
ChatGPT 18.1 (96.6%) [0] 56.2 (98.3%) [0] 59.2 (98.3%) [0] 14.1 (93.3%) [0] 64.7 (97.9%) [0] 66.3 (98.3%) [0] 34.7 (99.0%) [0] 73.2 [0] 73.7 (99.7%) [0]
GPT-4 35.1 (13.5%) [0] 59.6 (96.0%) [0] 61.2 (96.4%) [0] 33.8 (25.4%) [0] 65.8 (97.5%) [0] 68.4 (96.5%) [0] 54.4 (56.1%) [0] 76.4 (99.3%) [5] 78.8 (98.6%) [5]
Table 3: Best performance over few-shot settings of baselines for Vanilla prompt with searched results from Google. +5 and +10 represent different numbers of searched results appended in the inputs. We report in the form of F1-recall (answer rate) [best number of few shot examples] and omit the answer rate if it is 100%. Data with the highest F1-recall scores are marked in bold.
Model fast-changing slow-changing never-changing
+0 +5 +10 +0 +5 +10 +0 +5 +10
Intermlm-20B-Chat 15.0 [16] 46.8 [5] 49.5 [5] 18.1 [16] 59.3 [16] 61.7 [0] 34.7 [16] 67.2 [16] 68.9 [16]
Aquila2-34B-Chat 14.5 [16] 50.0 (99.6%) [0] 51.9 [0] 17.1 [16] 60.0 [0] 61.4 [0] 35.6 [5] 69.5 [0] 69.8 [0]
Yi-34B-Chat 23.2 [16] 54.7 [5] 57.4 [16] 30.4 [16] 67.8 [16] 68.5 [16] 47.0 [16] 75.3 [16] 77.3 [16]
Deepseek-67B-Chat 22.9 [16] 58.0 [16] 59.2 [16] 37.0 [5] 67.8 [16] 70.6 [16] 53.0 [16] 79.8 [16] 80.2 [16]
Qwen-72B-Chat 17.8 [16] 63.6 [5] 64.2 [16] 17.8 [16] 63.1 [0] 63.6 [5] 40.8 [16] 74.2 [16] 76.1 [16]
ChatGPT 17.9 (97.3%) [0] 57.6 (97.0%) [0] 61.4 (96.6%) [0] 13.9 (98.3%) [0] 65.5 (97.5%) [0] 65.7 (98.7%) [0] 36.0 (99.7%) [0] 75.2 (98.6%) [0] 74.9 (98.6%) [0]
GPT-4 22.1 (82.9%) [16] 67.3 (89.0%) [0] 68.0 (89.0%) [5] 19.8 (86.7%) [0] 73.9 (93.3%) [0] 74.7 (90.4%) [5] 48.2 (56.1%) [0] 81.6 (98.3%) [5] 83.5 (98.3%) [0]
Table 4: Best performance over few-shot settings of baselines for CoT prompt with searched results from Google. +5 and +10 represent different numbers of searched results appended in the inputs. We report in the form of F1-recall (answer rate) [best number of few shot examples] and omit the answer rate if it is 100%. Data with the highest F1-recall scores are marked in bold.
Model fast-changing slow-changing never-changing
+0 +5 +10 +0 +5 +10 +0 +5 +10
Intermlm-20B-Chat 16.9 [16] 49.5 [0] 52.4 [16] 20.0 [16] 61.2 [5] 63.3 [0] 34.6 [5] 68.2 [16] 71.3 [16]
Aquila2-34B-Chat 15.5 [16] 48.4 [16] 51.4 [0] 17.5 [16] 59.6 [16] 61.9 [0] 36.1 [16] 68.0 [16] 69.5 [0]
Yi-34B-Chat 22.8 [16] 54.6 [16] 57.0 [16] 30.6 [16] 68.2 [16] 68.5 [0] 47.7 [16] 77.1 [5] 76.8 [16]
Deepseek-67B-Chat 23.3 [5] 57.6 [16] 58.9 [5] 37.7 [5] 68.5 [16] 70.7 [16] 54.2 [16] 79.6 [5] 79.8 [5]
Qwen-72B-Chat 18.3 [16] 52.5 [16] 55.4 [16] 17.1 [0] 62.9 [16] 65.4 [16] 39.5 [16] 74.1 [16] 75.7 [16]
ChatGPT 19.2 [0] 58.9 [0] 61.7 [0] 15.9 [0] 65.8 (99.4%) [0] 67.6 (99.6%) [0] 35.6 [0] 75.2 (99.7%) [0] 76.5 (99.7%) [0]
GPT-4 29.1 (37.3%) [0] 61.8 (94.5%) [0] 64.9 (95.2%) [0] 28.3 (54.62%) [0] 71.6 (94.2%) [0] 71.6 (96.2%) [0] 54.9 (72.5%) [0] 80.7 (99.0%) [0] 82.6 (99.3%) [0]
Table 5: Best performance over few-shot settings of baselines for RaR prompt with searched results from Google. +5 and +10 represent different numbers of searched results appended in the inputs. We report in the form of F1-recall(answer rate)[best number of few shot examples] and omit the answer rate if it is 100%. Data with the highest F1-recall scores are marked in bold.
3.1 Experiment Setup
Baselines
We experiment with a series of baseline models pretrained with Chinese data, including representative OpenAI’s ChatGPT (gpt-3.5-turbo-1106) OpenAI (2022) and GPT-4 (gpt-4-1106-preview) OpenAI (2023), open-sourced Chinese-oriented models such as Internlm-20B-Chat InternLMTeam (2023), Aquila2-34B-Chat BAAI (2023), Yi-34B-Chat 01-ai (2023), Deepseek-67B-Chat DeepSeek-AI (2024), Qwen-72B-Chat Bai et al. (2023). In the close-book scenario, we only use the standalone LLM in answering questions with zero-shot and few-shot settings. For the open-book scenario, search engines are used for retrieving question-related results on the Internet and then fed into LLMs for reading.
Search Engines
Except for language models for information synthesis, we select three representative search engines to recall relevant passages from the Chinese Internet namely Google and Bing. These search engines are mainly used by Chinese people for information seeking. Baidu is omitted due to the difficulty in scraping its contents. The Top-5 and Top-10 searched results are provided to models in the open-book setting.
Prompt Design
Our prompt framework, which is in Chinese, could be framed as concatenation of $(\mathbf{i},\mathbf{d},\mathbf{e},\mathbf{q})$, in Figure 2 where $\mathbf{i}$ represents the instruction, $\mathbf{d}$ for question-answer pairs from crowdsourced labelers, $\mathbf{e}$ for search results and $\mathbf{q}$ for current question. Different instructions i are used with three widely adopted prompting styles, Vanilla, Chain-of-Thought (CoT) Wei et al. (2023) and Rephrase-and-Respond (RaR) Deng et al. (2023). Vanilla instruction is directly asking models to answer questions with the context. CoT instruction is asking models to first explain and analyze the question $\mathbf{q}$ step by step and then give their answers. RaR instruction, however, is asking models to first rephrase and expand the question $\mathbf{q}$ and then give their answers, which could be viewed as a complement of CoT Deng et al. (2023) as CoT is for diving deeper while RaR is for exploring broader. Besides, for demonstrations $\mathbf{d}$, we have used zero-shot and different few-shot settings, i.e., 5-shot and 16-shot. More specifically, our few-shot demonstrations are made up of human written questions and answers similar to CDQA dataset without contexts or other explanations as it costs longer time without any improvement.
3.2 Results and Analyses
Table 345 summarize best performances over few-shot prompting across different baselines for Vanilla, CoT and RaR prompts respectively. Our default search engine for analysis is Google as it is most widely used around the world.
Baseline Comparison
In Table 4, we see that GPT-4 achieves the best across all questions. Only GPT-4 reaches nearly or over 65, 70 and 80 in F1-recall for fast-changing, slow-changing and never-changing questions respectively with searched results which indicates that there is still room of improvement for other open-sourced models with Chinese instructions under the retrieval-augmented-generation scenario. Besides, the second-best model, Deepseek-67B-Chat, has shown great performance as it even surpasses GPT-4 on slow-changing and never-changing questions by Vanilla prompt with searched results in Table 3 and performs better (both in answer rate and F1-recall) than GPT-4 without searched results in slow-changing and never-changing questions especially in CoT prompt. This indicates that Deepseek-67B-Chat has stored more Chinese Internet knowledge as its alignment corpora mainly focus on Chinese and English while GPT-4 aligns with multilingual data. Moreover, it is worth noting that answer rates of ChatGPT and GPT-4 are often lower than 100% across all different prompts and question types especially for GPT-4 in close-book question answering which indicates that there are more hallucination reduction measures such as refusal of questions in GPT-4 than other models.
How do different styles of prompts work in LLMs?
As our evaluation scenarios are made up of close-book and open-book, different prompts just elicit different behaviors in LLMs. To rule out the influence of few-shot demonstrations, we use zero-shot setting in this analysis. We use GPT-4 and Deepseek models as closed and open-sourced representative models in the following analysis. In Figure 3, under close-book scenario, we see that Deepseek-67B-Chat has same answer rates over different prompts on different questions while there are different answering behaviors in GPT-4. Specifically, GPT-4 answers with great care in vanilla prompts with lowest answer rates but high F1-recall scores while GPT-4 suffers from hallucination in CoT and RaR prompts with at most +522% and +176% in answer rates but -43% and -17% in F1-recall scores compared to Vanilla prompt. For both models, Vanilla prompts outperform the other two kinds of prompts. This indicates that verbose explanation or expansion could increase hallucination especially when without evidence.
(a) Deepseek-67B-Chat
(b) GPT-4
Figure 3: F1-recall scores and Answer Rates of different prompts for LLMs in close-book scenario under zero-shot setting. We represent F1-recall scores with bar plots and answer rates with dotted lines.
(a) Deepseek-67B-Chat
(b) GPT-4
Figure 4: F1-recall scores and Answer Rates of different prompts for LLMs in open-book scenario under zero-shot setting. We represent F1-recall scores with bar plots and answer rates with dotted lines.
In Figure 4, under open-book scenario [^3] We use Top-10 searched results as more searched results generally improve the results of LLMs from Table 345., we see that Deepseek-67B-Chat still shares same answer rates across different prompts and question types and vanilla prompt is the best for it, which indicates that CoT and RaR take risks in inducing more hallucinated answers. For GPT-4, answer rates have less gaps between different prompts (all over 90%) and F1-recall scores are all increasing dramatically compared to close-book counterparts, representing adding contextual information elicits LLMs in answering questions more efficiently. Besides, with search results, CoT and RaR both outperform Vanilla prompt and CoT performs the best in GPT-4 with less hallucination, i.e., lower answer rate, and higher F1-recall score. This indicates that CoT and RaR could improve LLMs on complex tasks but CoT elicits more reasoning abilities to directly improve the answering.
Nevertheless, model sizes and training data are both fundamental for these prompts to work. In Figure 5, not every model improves with CoT or RaR compared to Vanilla prompt. For example, Deepseek-34B-Chat’s and Qwen-72B-Chat’s performance decrease in CoT and RaR; ChatGPT and Internlm-20B-Chat prefer RaR while GPT-4 and Yi-34B-Chat prefer CoT for larger gains; Aquila2-34B-Chat is robust to all prompt types.
Figure 5: F1-recall scores averaged over all three different questions for all models with different prompts in open-book scenario under zero-shot setting. We present F1-recall score only since all answer rates $\geq$ 90%.
Does few-shot prompting always work for all LLMs?
For better analyzing the influence of few-shot prompting, we collect close-book and open-book results of all LLMs in zero-shot setting and vanilla prompt. In Figure 6, based on nearly 100% answer rate, five in close-book scenario and four in open-book scenario out of all five open-sourced models have better performance with more few-shot demonstrations, which are sampled in the same data distribution during generating CDQA.
(a) Close-book
(b) Open-book (Top-10 searched results)
Figure 6: F1-recall scores averaged over all types of questions for different models with different few-shot settings in close-book and open-book scenarios.
Models close-book open-book
0-shot 5-shot 16-shot 0-shot 5-shot 16-shot
ChatGPT 96.3 95.0 96.7 98.8 99.7 99.9
GPT-4 31.7 52.2 64.7 97.4 97.7 98.0
Table 6: Answer rates for ChatGPT and GPT-4 averaged on all types of questions in close-book and open-book scenarios with different few-shot settings.
However, we also notice that ChatGPT and GPT-4 have shown different trends compared to open-sourced models, i.e., more few-shot examples lead to decreases in F1-recall scores. Therefore, we check the answer rates of ChatGPT and GPT-4 in Table 6 where ChatGPT stays in fairly high answer rates ($\geq 95\%$) and GPT-4 increases its answer rates with more few-shot examples. Combined with their monotonic decrease in F1-recall scores, we show that ChatGPT and GPT-4 hallucinate more with more few-shot examples in CDQA. This indicates the challenge of CDQA and different LLMs with different capabilities struggle in different ways.
How do different search engines help?
For a more detailed comparison between search engines across all baselines, we use vanilla prompt since CoT performs better in LLMs and zero-shot setting with Top-10 searched results from different search engines. In Figure 7, searched results from Google consistently outperforms Bing among all baseline models, which indicates that the Google currently provides the most useful retrieved evidence for question answering about Chinese news.
Figure 7: F1-recall scores averaged over all questions for different models with different search engines.
How do LLMs perform across different answer types?
As answers in CDQA are mainly entities from news, we conduct analysis across different answer types for two representative open-sourced and closed LLMs, i.e., Deepseek-67B-Chat and GPT-4. In Figure 8, we observe that GPT-4’s internal knowledge is poorer than Chinese-oriented models such as Deepseek-67B-Chat for Chinese users. However, with enough retrieved evidence, GPT-4 has stronger abilities in learning from contexts than Deepseek-67B-Chat where this “learning efficiency”, i.e., the ratio of gaps between open-book scores and close-book scores to the close-book could reach at most 1370% compared to 219% in Deepseek-67B-Chat.
Figure 8: F1-recall scores on different answer types for Deepseek-67B-Chat and GPT-4 in close-book and open-book scenarios with Vanilla prompt. We use Top 10 searched results from Google. Deepseek-67B-Chat holds 100% answer rate in all time while GPT-4 drastically increase its answer rate to 100% with searched results from Google.
Moreover, from Figure 8, we also could notice that the “quantity” group is hardest for models to figure out the correct answers, which is due to the granularity of answers and the need of reasoning abilities.
4 Related Work
Temporal and Dynamic QA Benchmark
StreamingQA Liska et al. (2022) is a QA dataset where questions are human writen or LM-generated on given dates, showing how open-book and close-book QA models adapt to new knowledge over time and importance of retrieval augmentation in up-to-date search space. TimeQA Chen et al. (2021) are formed from extracted evolving facts in WikiData by manual extraction and verification while we extract entities to directly formulate them as answer candidates based on the documents. RealTimeQA Kasai et al. (2022), a dynamic QA benchmark with automatic weekly updates from the weekly News Quiz section in social media such as CNN, is most related to our semi-automatic question generation with latest Chinese news corpus.
English QA benchmark
Question answering is a long-standing task in NLP area Wang et al. (2024); Li et al. (2024), ranging from classic single-turn benchmarks such as SQuAD Rajpurkar et al. (2016, 2018), TriviaQA Joshi et al. (2017) and Natural Questions Kwiatkowski et al. (2019) to conversational QA like TopiOCQA Adlakha et al. (2022). Also, it could also be categorized by normal human intelligence and human experts intelligence. For example, CommonsenseQA Talmor et al. (2019) comes up with questions from ConceptNet in discriminating the target concepts with commonsense while GPQA Rein et al. (2023) consists of questions generated from graduate-level students. Although these benchmarks have provided efficient and effective evaluation metrics and covered a variety of topics, they are mostly static with a focus on less-evolving knowledge, which might already be or going to be included in the pre-training data of large language models.
Chinese QA benchmark
In contrast to prosperous English QA benchmarks, Chinese counterparts are still under-explored. DuReader He et al. (2018) is a classic free-form QA benchmark collected by Baidu from its own products and CLUE Xu et al. (2020) is the first large scale NLU benchmark in Chinese. After the recent debut of powerful large language models, a series of Chinese QA benchmarks are proposed for better evaluating them. C-Eval Huang et al. (2023b) is a multiple-choice questions answering dataset from Chinese Standard Exams. WebCPM Qin et al. (2023) collects questions from web forums through web searching and browsing and SuperCLUE Xu et al. (2023) is a comprehensive Chinese benchmark for question answering in aligning users needs. But they all suffer from either data leakage or the risk of saturated performance which hinders the accurate evaluation on questions requiring fresh knowledge to answer as static questions are readily overfitted.
5 Conclusion
The creation of CDQA addresses the urgent need for the evaluation of Chinese LLMs, thereby improving LLM-driven applications for Chinese users. Given the cultural influences in LLMs’ training data, it is our aspiration that CDQA will foster development in various capabilities of LLMs, particularly within Chinese contexts. While CDQA progresses further with a semi-automatic generation pipeline with more data than FreshQA, we acknowledge that it is far from a perfect LLM evaluation. Other critical dimensions, including tool learning, LLMs safety, and robustness, remain to be explored. However, we believe that our constructed CDQA and the series of insights obtained based on it will provide valuable resources and guidance for subsequent research on Chinese LLMs.
Limitations
One of the limitations of our work is that the language we study is Chinese only. As the two most widely used languages in the world, English and Chinese have always been equally valued and widely concerned in the NLP community. In fact, our work is inspired by previous FreshQA in the English scenario and aims to provide similar data resources to Chinese LLMs researchers. We also encourage and welcome more researchers from other languages to engage in similar research.
In addition, another limitation that cannot be ignored is how to keep our CDQA updated. Because CDQA focuses on questions whose answers change dynamically, it is critical to ensure that the answers to questions in CDQA are always correct and up-to-date. Therefore, we also commit to updating our CDQA regularly and providing researchers with the latest version of CDQA for LLMs evaluation.
Ethics Statement
The task we focus on is the evaluation of LLMs, and the LLMs we evaluate are all public and widely used LLMs, so they do not bring potential ethical risks. The data samples of CDQA that we collect have been manually cleaned and pre-processed to ensure that they do not contain any data that will cause moral risks, such as politically sensitive, violent, and private data. In addition, we also have signed legal labor contracts with the human annotators we employ, and pay them higher than market prices based on their workload.
Appendix A Tag Taxonomy of CDQA
The tag taxonomy of CDQA is presented in Table 7.
Category Description Example
fast-changing
The answer to the question is prone to changing within one year
(How many sessions has the Maritime Silk Road Cultural Heritage Forum been held?, Four)
slow-changing
The answer to the question is prone to changing in several years
(Which ancient city site in China has recently been recognized as a UNESCO World Cultural Heritage?; the Liangzhu Ancient City Site)
never-changing
The answer to the question is from static knowledge such as scientific theories, historical facts and so on
(In rural areas during winter heating, it is necessary to guard against the risk of poisoning from which gas?; Carbon monoxide)
  person
Specific individual, usually referring to a human being.
(Who among the current representatives of the Fuxin County People’s Congress was one of the first batch of anti-epidemic heroes to rush to support Wuhan?; Xin Li)
location
Geographical position.
(Which province has recently strengthened the regulation of the intellectual property agency industry?; Hainan)
time
Points or intervals of a continuous sequence of events or conditions.
(In which year was the recent "Haikou Cup" sailing competition held?; 2023)
event
Something that happens, which can be planned or spontaneous.
(What themed event was recently launched in Suzhou High-speed Railway New Town to promote the development of private enterprises?; "Suzhou Sentiments, Private Enterprises Connected at Heart")
artificial work
Items or intellectual achievements created by humans, which have artistic, academic, or practical value.
(What is the latest TV series aired starring Xin Jiang?; As Long As We Are Together)
group
Entities formed by multiple individuals for a specific purpose.
(Which undergraduate university is rencently established in the Ningxia Hui Autonomous Region recently?; Ningxia Minjiang Institute of Applied Technology)
nature
Phenomena or entities in the natural world.
(Please explain to me what Nucleases is?; Small RNA molecules with catalytic function, belonging to the category of biological catalysts?; capable of degrading specific mRNA sequences.)
quantity
Numeric value for times or stuff.
(How many base pairs in human Y chromosome have been observed from the latest sequencing results?;More than 30 million)
other
Other answer not classified to the above categories.
(Is there any fee for withdrawing WeChat balance to bank card?; Yes)
Table 7: Descriptions and examples of question tags (first three rows) and answer types (last nine rows). We represent (<question>; <answer>) as examples. Original language for these examples is Chinese. We translate them here for better preview.
Appendix B Dataset Distributions
Knowledge types for queries and answer types are visualized in the following Figure 910. More specifically, we have further visualized the answer type distributions in each question tag. From Figure 11, Figure 12 and Figure 13, we see that nearly 80% of slow changing questions are about person and group. Except for person and group, artificial work should be the third largest category for answers, which includes jobs, titles, knowledge and so on. These observations are all consistent with our data sources as information for the protagonists, places and events are compulsory and most frequent in news reports. Besides, percentages of time reach the maximum in never-changing tag as currently most of questions answered with time are about the frequencies.
Figure 9: Distributions of question tags for full data.
Figure 10: Distribution of answer types for full data.
Figure 11: Distributions of answer types for fast-changing questions.
Figure 12: Distributions of answer types for slow-changing questions.
Figure 13: Distributions of answer types for never-changing questions.
Appendix C Translated Chinese Prompts
The translated prompt framework is illustrated in Figure 14.
Figure 14: The Chinese prompt framework for Figure 2.

### 02 Mar 2024 v2

@@
 Extensive experiments and valuable insights suggest that our proposed CDQA is challenging and worthy of more further study.
-We believe that the benchmark we provide will become the key data resource for improving LLMs’ Chinese question-answering ability in the future.
+We believe that the benchmark we provide will become one of the key data resources for improving LLMs’ Chinese question-answering ability in the future.
 1 Introduction
@@
 Deepseek-67B-Chat 24.3 [0] 57.5 [16] 58.4 [16] 37.2 [16] 68.3 [16] 70.0 [16] 53.1 [0] 78.8 [16] 79.2 [5]
-Qwen-72B-Chat 18.6 [16] 51.7 [16] 54.7 [16] 16.9 [16] 63.1 [0] 64.6 [16] 41.5 [16] 74.6 [16] 75.4 [16]
 ChatGPT 18.1 (96.6%) [0] 56.2 (98.3%) [0] 59.2 (98.3%) [0] 14.1 (93.3%) [0] 64.7 (97.9%) [0] 66.3 (98.3%) [0] 34.7 (99.0%) [0] 73.2 [0] 73.7 (99.7%) [0]
@@
 Deepseek-67B-Chat 22.9 [16] 58.0 [16] 59.2 [16] 37.0 [5] 67.8 [16] 70.6 [16] 53.0 [16] 79.8 [16] 80.2 [16]
-Qwen-72B-Chat 17.8 [16] 63.6 [5] 64.2 [16] 17.8 [16] 63.1 [0] 63.6 [5] 40.8 [16] 74.2 [16] 76.1 [16]
 ChatGPT 17.9 (97.3%) [0] 57.6 (97.0%) [0] 61.4 (96.6%) [0] 13.9 (98.3%) [0] 65.5 (97.5%) [0] 65.7 (98.7%) [0] 36.0 (99.7%) [0] 75.2 (98.6%) [0] 74.9 (98.6%) [0]
@@
 Deepseek-67B-Chat 23.3 [5] 57.6 [16] 58.9 [5] 37.7 [5] 68.5 [16] 70.7 [16] 54.2 [16] 79.6 [5] 79.8 [5]
-Qwen-72B-Chat 18.3 [16] 52.5 [16] 55.4 [16] 17.1 [0] 62.9 [16] 65.4 [16] 39.5 [16] 74.1 [16] 75.7 [16]
 ChatGPT 19.2 [0] 58.9 [0] 61.7 [0] 15.9 [0] 65.8 (99.4%) [0] 67.6 (99.6%) [0] 35.6 [0] 75.2 (99.7%) [0] 76.5 (99.7%) [0]
@@
 Baselines
-We experiment with a series of baseline models pretrained with Chinese data, including representative OpenAI’s ChatGPT (gpt-3.5-turbo-1106) OpenAI (2022) and GPT-4 (gpt-4-1106-preview) OpenAI (2023), open-sourced Chinese-oriented models such as Internlm-20B-Chat InternLMTeam (2023), Aquila2-34B-Chat BAAI (2023), Yi-34B-Chat 01-ai (2023), Deepseek-67B-Chat DeepSeek-AI (2024), Qwen-72B-Chat Bai et al.
-(2023).
+We experiment with a series of baseline models pretrained with Chinese data, including representative OpenAI’s ChatGPT (gpt-3.5-turbo-1106) OpenAI (2022) and GPT-4 (gpt-4-1106-preview) OpenAI (2023), open-sourced Chinese-oriented models such as Internlm-20B-Chat InternLMTeam (2023), Aquila2-34B-Chat BAAI (2023), Yi-34B-Chat 01-ai (2023), Deepseek-67B-Chat DeepSeek-AI (2024).
 In the close-book scenario, we only use the standalone LLM in answering questions with zero-shot and few-shot settings.
@@
 In Figure 5, not every model improves with CoT or RaR compared to Vanilla prompt.
-For example, Deepseek-34B-Chat’s and Qwen-72B-Chat’s performance decrease in CoT and RaR; ChatGPT and Internlm-20B-Chat prefer RaR while GPT-4 and Yi-34B-Chat prefer CoT for larger gains; Aquila2-34B-Chat is robust to all prompt types.
+For example, Deepseek-34B-Chat’s performance decrease in CoT and RaR; ChatGPT and Internlm-20B-Chat prefer RaR while GPT-4 and Yi-34B-Chat prefer CoT for larger gains; Aquila2-34B-Chat is robust to all prompt types.
 Figure 5: F1-recall scores averaged over all three different questions for all models with different prompts in open-book scenario under zero-shot setting.
@@
 For a more detailed comparison between search engines across all baselines, we use vanilla prompt since CoT performs better in LLMs and zero-shot setting with Top-10 searched results from different search engines.
-In Figure 7, searched results from Google consistently outperforms Bing among all baseline models, which indicates that the Google currently provides the most useful retrieved evidence for question answering about Chinese news.
+In Figure 7, searched results from Google slightly outperforms Bing among all baseline models, which indicates that the Google currently provides the most useful retrieved evidence for question answering about Chinese news.
 Figure 7: F1-recall scores averaged over all questions for different models with different search engines.
@@
 However, we believe that our constructed CDQA and the series of insights obtained based on it will provide valuable resources and guidance for subsequent research on Chinese LLMs.
+In the future, we will conduct more in-depth analyses of the capabilities of LLMs based on CDQA and investigate how to enhance the LLMs’ ability to handle dynamic questions.
+This will empower LLMs to better cope with the complex and ever-changing real-world application environments.
 Limitations